# Classification of $4$-dimensional complex Poisson algebras

### v1

@@ Abstract.
The present paper is devoted to the complete classification of $4$-dimensional complex Poisson algebras, taking into account a classification, up to isomorphism, of the complex commutative associative algebras of dimension $4$, as well as by using a Lie algebra classification, up to isomorphism.
@@
The second author was supported by the Centre for Mathematics of the University of Coimbra - UIDB/00324/2020, funded by the Portuguese Government through FCT/MCTES. Third author is supported by the PCI of the UCA ‘Teoría de Lie y Teoría de Espacios de Banach’ and by the PAI with project number FQM298.
@@ 1. Introduction
Poisson algebras are vector spaces equipped with two binary operations: a commutative product and a Lie bracket, which satisfy a certain compatibility condition known as the Leibniz rule. Introduced in Hamiltonian mechanics as the dual of the category of classical mechanical systems, Poisson algebras have a fundamental role in the study of quantum groups, differential geometry, noncommutative geometry, integrable systems, quantum field theory or vertex operator algebras (see [4, 5, 6, 7, 8, 9, 10]). Additionally, this class of algebras plays an important role in many areas of mathematics including symplectic geometry, representation theory, quantum field theory and algebraic geometry. Poisson algebras can be thought of as the algebraic counterpart of Poisson manifolds which are smooth manifolds $M$ whose commutative algebra $C^{\infty}(M,\,{\mathbb{R}})$ of real smooth functions is endowed with a Lie bracket [,][-,-][ - , - ] satisfying the Leibniz rule, i.e., $C^{\infty}(M,\,{\mathbb{R}})$ is a Poisson algebra. Here, they are used to study geometric structures that preserve certain properties under deformation.
@@
In [1] the authors developed a method to obtain the algebraic classification of Poisson algebras defined on a commutative associative algebra, and they applied it to obtain the classification of the $3$-dimensional complex Poisson algebras. In addition, they also study the algebraic classification of the Poisson algebras defined on a commutative associative null-filiform or filiform algebra. This method focus on presenting a procedure to classify all the Poisson algebras associated with a given commutative associative algebra. We briefly explain the method. Pick an arbitrary commutative associative algebra $\mathcal{P}$. Further, compute the set $Z^{2}(\mathcal{P},\mathcal{P})$ of all skew symmetric bilinear maps on $\mathcal{P}$ satisfying some adequate conditions (Definition 2.3). It is proved that for any Poisson algebra $(\mathcal{P},\cdot,[-,-])$ there exists $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$ such that the algebra is isomorphic to a Poisson algebra $(\mathcal{P},\cdot,[-,-]_{\theta})$ associated with the given commutative associative algebra $\mathcal{P}$. Find the orbits of the automorphisms group ${\rm Aut}(\mathcal{P})$ on $Z^{2}(\mathcal{P},\mathcal{P})$ by a proper action (3). It turns out that, by choosing a representative $\theta$ from each orbit, it is obtained all Poisson algebras $(\mathcal{P},\cdot,[-,-]_{\theta})$, up to isomorphism. So it is obtained all the Poisson algebras associated with a given commutative associative algebra.
In the present paper, the authors produce a complete classification of $4$-dimensional complex Poisson algebras, by applying the above results, taking into account the classification, up to isomorphism, of the complex commutative associative algebras of dimension $4$ presented in [2], as well as, using the Lie algebra classification, up to isomorphism, given in [3].
@@ A Poisson algebra is a vector space $\mathcal{P}$ endowed with two bilinear operations:
    An commutative associative multiplication denoted by $-\cdot-:\mathcal{P}\times\mathcal{P}\to\mathcal{P}$;
@@
    A Lie algebra multiplication denoted by $[-,-]:\mathcal{P}\times\mathcal{P}\to\mathcal{P}$.
@@
(1) $[x\cdot y,z]=[x,z]\cdot y+x\cdot[y,z],$
@@
Observe that, for $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$, we may define on $\mathcal{P}$ a bracket $[-,-]_{\theta}:\mathcal{P}\times\mathcal{P}\to\mathcal{P}$ by
(2) $[x,y]_{\theta}:=\theta(x,y),$
@@
Similarly, we can construct an analogous method for classifying the $4$-dimensional Poisson algebras from the classification of Lie algebras of dimension four.
@@
[…]
Now, if $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$, then $\theta$ can be uniquely written as $\theta(x,y)=\sum_{i=1}^{n}B_{i}(x,y)e_{i}$ where $B_{1},B_{2},\dots,B_{n}$ is a sequence of skew symmetric bilinear forms on $\mathcal{P}$. Also, we may write $\theta=(B_{1},B_{2},\dots,B_{n})$ Let $\phi^{-1}\in{\rm Aut}(\mathcal{P})$ be given by the matrix $(b_{ij})$. If $(\theta*\phi)(x,y)=\sum_{i=1}^{n}B_{i}^{\prime}(x,y)e_{i}$ then $B_{i}^{\prime}=\sum_{j=1}^{n}b_{ij}\phi^{t}B_{j}\phi$.
Remark 2.2.
@@
3. Poisson algebras of dimension $4$
From now, we present the classification, up to isomorphism, of the Poisson algebras for dimension $4$ over the field of complex numbers $\mathbb{C}$. For simplicity, every time we write the multiplication table of a Poisson algebra the products of basic elements whose values are zero or can be recovered by the commutativity, in the case of $\cdot$-- ⋅ -, or by the anticommutativity, in the case of [,][-,-][ - , - ], are omitted. First we recall the classification, up to isomorphism, of the complex commutative associative algebras of dimension $4$ presented in [2]:
@@ 3. Poisson algebras of dimension $4$
[2] Let $\mathcal{A}$ be a complex commutative associative algebra of dimension $4$. Then $\mathcal{A}$ is isomorphic to one of the following algebras:
@@
Let $\mathcal{P}$ be a complex Poisson algebra of dimension $4$. Then $\mathcal{P}$ is isomorphic to one of the Lie algebras given in [3, Lemma 3] or to one of the following algebras:
@@
  • $\mathcal{P}_{4,2}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{3}\right]=e_{2}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,3}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{2}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,4}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{3}\right]=e_{3},\left[e_{1},e_{4}\right]=\alpha e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,5}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{4}\right]=e_{2},\left[e_{1},e_{3}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,6}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{2},\left[e_{1},e_{3}\right]=e_{3},\left[e_{1},e_{4}\right]=-e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,7}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{3}\right]=e_{3},\left[e_{1},e_{4}\right]=e_{3}+e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,8}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,9}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{3}\right]=e_{2},\left[e_{1},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,10}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{2},\left[e_{1},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,11}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,12}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},$\\ $\left[e_{1},e_{4}\right]=e_{2},\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,13}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=\alpha e_{3},\left[e_{1},e_{4}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,14}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,15}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{3},\left[e_{1},e_{2}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,16}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=\alpha e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,17}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,18}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{3},\left[e_{2},e_{4}\right]=ie_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,19}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{3},\left[e_{2},e_{4}\right]=ie_{3},\left[e_{1},e_{2}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,20}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=\alpha e_{3},\left[e_{1},e_{4}\right]=e_{4},\left[e_{2},e_{4}\right]=ie_{4}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,22}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},e_{1}\cdot e_{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,23}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},e_{1}\cdot e_{2}=e_{3},$\\ $\left[e_{1},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,25}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{3},e_{2}^{2}=e_{4},$\\ $\left[e_{1},e_{2}\right]=e_{3}+\alpha e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,26}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=e_{1},\left[e_{1},e_{3}\right]=2e_{4},\left[e_{2},e_{4}\right]=-e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,27}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,28}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{2}^{2}=e_{3},$\\ $\left[e_{1},e_{2}\right]=\alpha e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,29}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{3}^{2}=e_{4},$\\ $\left[e_{1},e_{3}\right]=-e_{1},\left[e_{2},e_{3}\right]=e_{2},\left[e_{1},e_{2}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,30}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{3}^{2}=e_{4},$\\ $\left[e_{1},e_{3}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,31}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}\cdot e_{2}=e_{4},e_{3}^{2}=e_{4},$\\ $\left[e_{1},e_{2}\right]=\alpha e_{4}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,33}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{2},e_{1}\cdot e_{2}=e_{4},e_{3}^{2}=e_{4},$\\ $\left[e_{1},e_{3}\right]=e_{4}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,39}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{2}^{2}=e_{2},e_{2}\cdot e_{3}=e_{3},e_{2}\cdot e_{4}=e_{4},$\\ $\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,41}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},$\\ $\left[e_{2},e_{4}\right]=\alpha e_{2},\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,42}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},$\\ $\left[e_{2},e_{4}\right]=e_{2},\left[e_{3},e_{4}\right]=e_{2}+e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,44}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},$\\ $\left[e_{3},e_{4}\right]=e_{2}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,45}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},$\\ $\left[e_{3},e_{4}\right]=e_{2},\left[e_{2},e_{3}\right]=e_{3},\left[e_{2},e_{4}\right]=-e_{4}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,47}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},e_{2}^{2}=e_{3},$\\ $\left[e_{2},e_{4}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,48}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},e_{2}^{2}=e_{3},$\\ $\left[e_{2},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,50}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},e_{1}\cdot e_{4}=e_{4},e_{2}^{2}=e_{4},e_{3}^{2}=e_{4},$\\ $\left[e_{2},e_{3}\right]=\alpha e_{4}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,54}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},e_{1}\cdot e_{3}=e_{3},$\\ $\left[e_{2},e_{3}\right]=e_{2}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,57}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{2}^{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,60}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{1}\cdot e_{2}=e_{2},$\\ $\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,62}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},$\\ $\left[e_{2},e_{4}\right]=\alpha e_{2},\left[e_{3},e_{4}\right]=e_{3}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,63}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},$\\ $\left[e_{2},e_{4}\right]=e_{2},\left[e_{3},e_{4}\right]=e_{2}+e_{3}.$\end{tabular}\right.$
@@
  • $\mathcal{P}_{4,65}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},$\\ $\left[e_{3},e_{4}\right]=e_{2}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,66}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},$\\ $\left[e_{3},e_{4}\right]=e_{2},\left[e_{2},e_{3}\right]=e_{3},\left[e_{2},e_{4}\right]=-e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,67}^{\alpha}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{2}\cdot e_{3}=e_{4},$\\ $\left[e_{2},e_{3}\right]=\alpha e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,68}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{2}^{2}=e_{3},$\\ $\left[e_{2},e_{4}\right]=e_{4}.$\end{tabular}\right.$
  • $\mathcal{P}_{4,69}:\left\{\begin{tabular}[]{l}$e_{1}^{2}=e_{1},e_{2}^{2}=e_{3},$\\ $\left[e_{2},e_{4}\right]=e_{3}.$\end{tabular}\right.$
@@
      $\frac{1}{\alpha_{2}^{2}\alpha_{4}^{2}}\allowbreak\begin{pmatrix}\alpha_{2}\alpha_{4}&0&0&0\\ 0&\alpha_{2}\alpha_{4}&0&0\\ 0&0&1&\alpha_{1}\\ 0&0&0&\alpha_{4}\end{pmatrix},\frac{1}{\alpha_{2}^{2}\alpha_{4}^{2}}\allowbreak\begin{pmatrix}-\alpha_{2}\alpha_{4}&0&0&0\\ 0&\alpha_{2}\alpha_{4}&0&0\\ 0&0&1&-\alpha_{1}\\ 0&0&0&-\alpha_{4}\end{pmatrix}.\allowbreak\allowbreak$
@@
  • $\alpha_{1}=\alpha_{2}=0$. Then $\theta\ast\phi=\theta$ for any ϕitalic-$\phi\in$Aut$\left(\mathcal{A}_{06}\right)$. Thus we have the representatives $\theta^{\alpha}=\left(0,0,0,\alpha\Delta_{1,2}\right)$ and $\theta^{\alpha},$ $\theta^{\beta}$ are in the same orbit if and only if $\alpha=\beta$. Therefore, we obtain the Poisson algebras $\mathcal{P}_{4,28}^{\alpha}$.
@@
    $\alpha_{2}-\alpha_{6}\alpha$ $=$ $0$
@@
    Then $\theta\ast\phi=\left(0,\alpha\Delta_{2,4},\Delta_{3,4},0\right)$ So we obtain the Poisson algebras $\mathcal{P}_{4,41}^{\alpha}$. If $\lambda\neq 0$, then $\alpha=-1$. Further if we choose $\phi$ as follows:
@@
for some $\alpha_{1},\ldots,\alpha_{9}\in\mathbb{C}$. Since Aut$\left(\mathcal{A}_{15}\right)=$Aut$\left(\mathcal{A}_{27}\right)$ and $Z^{2}\left(\mathcal{A}_{15},\mathcal{A}_{15}\right)=Z^{2}\left(\mathcal{A}_{27},\mathcal{A}_{27}\right)$, we obtain the algebras $\mathcal{P}_{4,62}^{\alpha},$ $\mathcal{P}_{4,63},$ $\mathcal{P}_{4,64},\mathcal{P}_{4,65},\mathcal{P}_{4,66}$.

### v2

# Classification of 44-dimensional complex Poisson algebras

@@ Abstract.
The present paper is devoted to the complete classification of 44-dimensional complex Poisson algebras, taking into account a classification, up to isomorphism, of the complex commutative associative algebras of dimension 44, as well as by using a Lie algebra classification, up to isomorphism.
@@
The second author is supported by the PCI of the UCA ‘Teoría de Lie y Teoría de Espacios de Banach’ and by the PAI with project number FQM298.
@@ 1. Introduction
Poisson algebras are vector spaces equipped with two binary operations: a commutative product and a Lie bracket, which satisfy a certain compatibility condition known as the Leibniz rule. Introduced in Hamiltonian mechanics as the dual of the category of classical mechanical systems, Poisson algebras have a fundamental role in the study of quantum groups, differential geometry, noncommutative geometry, integrable systems, quantum field theory or vertex operator algebras (see [4, 5, 6, 7, 8, 9, 10]). Additionally, this class of algebras plays an important role in many areas of mathematics including symplectic geometry, representation theory, quantum field theory and algebraic geometry. Poisson algebras can be thought of as the algebraic counterpart of Poisson manifolds which are smooth manifolds $M$ whose commutative algebra $C^{\infty}(M,\,{\mathbb{R}})$ of real smooth functions is endowed with a Lie bracket [,][-,-] satisfying the Leibniz rule, i.e., $C^{\infty}(M,\,{\mathbb{R}})$ is a Poisson algebra. Here, they are used to study geometric structures that preserve certain properties under deformation.
@@
In [1] the authors developed a method to obtain the algebraic classification of Poisson algebras defined on a commutative associative algebra, and they applied it to obtain the classification of the 33-dimensional complex Poisson algebras. In addition, they also study the algebraic classification of the Poisson algebras defined on a commutative associative null-filiform or filiform algebra. This method focus on presenting a procedure to classify all the Poisson algebras associated with a given commutative associative algebra. We briefly explain the method. Pick an arbitrary commutative associative algebra $\mathcal{P}$. Further, compute the set $Z^{2}(\mathcal{P},\mathcal{P})$ of all skew symmetric bilinear maps on $\mathcal{P}$ satisfying some adequate conditions (Definition 2.3). It is proved that for any Poisson algebra $(\mathcal{P},\cdot,[-,-])$ there exists $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$ such that the algebra is isomorphic to a Poisson algebra $(\mathcal{P},\cdot,[-,-]_{\theta})$ associated with the given commutative associative algebra $\mathcal{P}$. Find the orbits of the automorphisms group ${\rm Aut}(\mathcal{P})$ on $Z^{2}(\mathcal{P},\mathcal{P})$ by a proper action (3). It turns out that, by choosing a representative $\theta$ from each orbit, it is obtained all Poisson algebras $(\mathcal{P},\cdot,[-,-]_{\theta})$, up to isomorphism. So it is obtained all the Poisson algebras associated with a given commutative associative algebra.
In the present paper, the authors produce a complete classification of 44-dimensional complex Poisson algebras, by applying the above results, taking into account the classification, up to isomorphism, of the complex commutative associative algebras of dimension 44 presented in [2], as well as, using the Lie algebra classification, up to isomorphism, given in [3].
@@ A Poisson algebra is a vector space $\mathcal{P}$ endowed with two bilinear operations:
    An commutative associative multiplication denoted by $\cdot-:\mathcal{P}\times\mathcal{P}\to\mathcal{P}$;
@@
    A Lie algebra multiplication denoted by $\mathcal{P}\times\mathcal{P}\to\mathcal{P}$.
@@
(1) $x\cdot y,z]=[x,z]\cdot y+x\cdot[y,z],$
@@
Observe that, for $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$, we may define on $\mathcal{P}$ a bracket $_{\theta}:\mathcal{P}\times\mathcal{P}\to\mathcal{P}$ by
(2) $x,y]_{\theta}:=\theta(x,y),$
@@
Similarly, we can construct an analogous method for classifying the 44-dimensional Poisson algebras from the classification of Lie algebras of dimension four.
@@
$\Delta_{i,j}(e_{l},e_{m}):=\left\{\begin{tabular}[]{ll}$1,$&if $(i,j)=(l,m),$\\ $-1,$&if $(i,j)=(m,l),$\\ $0,$&otherwise.\end{tabular}\right$.
Now, if $\theta\in Z^{2}(\mathcal{P},\mathcal{P})$, then $\theta$ can be uniquely written as $\theta(x,y)=\sum_{i=1}^{n}B_{i}(x,y)e_{i}$ where $B_{1},B_{2},\dots,B_{n}$ is a sequence of skew symmetric bilinear forms on $\mathcal{P}$. Also, we may write $\theta=(B_{1},B_{2},\dots,B_{n})$ . Let $\phi^{-1}\in{\rm Aut}(\mathcal{P})$ be given by the matrix $(b_{ij})$. If $(\theta*\phi)(x,y)=\sum_{i=1}^{n}B_{i}^{\prime}(x,y)e_{i}$ then $B_{i}^{\prime}=\sum_{j=1}^{n}b_{ij}\phi^{t}B_{j}\phi$.
Remark $2.$
@@
3. Poisson algebras of dimension 44
From now, we present the classification, up to isomorphism, of the Poisson algebras for dimension 44 over the field of complex numbers $\mathbb{C}$. For simplicity, every time we write the multiplication table of a Poisson algebra the products of basic elements whose values are zero or can be recovered by the commutativity, in the case of -$\cdot$-, or by the anticommutativity, in the case of [,][-,-], are omitted. First we recall the classification, up to isomorphism, of the complex commutative associative algebras of dimension 44 presented in [2]:
@@ 3. Poisson algebras of dimension 44
[2] Let $\mathcal{A}$ be a complex commutative associative algebra of dimension 44. Then $\mathcal{A}$ is isomorphic to one of the following algebras:
@@
Let $\mathcal{P}$ be a complex Poisson algebra of dimension 44. Then $\mathcal{P}$ is isomorphic to one of the Lie algebras given in [3, Lemma 3] or to one of the following algebras:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
  • $\alpha_{1}=\alpha_{2}=0$. Then $\theta\ast\phi=\theta$ for any $\phi\inAut$$\left(\mathcal{A}_{06}\right)$. Thus we have the representatives $\theta^{\alpha}=\left(0,0,0,\alpha\Delta_{1,2}\right)$ and $\theta^{\alpha},$ $\theta^{\beta}$ are in the same orbit if and only if $\alpha=\beta$. Therefore, we obtain the Poisson algebras $\mathcal{P}_{4,28}^{\alpha}$.
@@
    $\alpha_{2}-\alpha_{6}\alpha$ $=$ $0.$
@@
    Then $\theta\ast\phi=\left(0,\alpha\Delta_{2,4},\Delta_{3,4},0\right)$ . So we obtain the Poisson algebras $\mathcal{P}_{4,41}^{\alpha}$. If $\lambda\neq 0$, then $\alpha=-1$. Further if we choose $\phi$ as follows:
@@
for some $\alpha_{1},\ldots,\alpha_{9}\in\mathbb{C}$. Since $\left(\mathcal{A}_{15}\right)=Aut$$\left(\mathcal{A}_{27}\right)$ and $Z^{2}\left(\mathcal{A}_{15},\mathcal{A}_{15}\right)=Z^{2}\left(\mathcal{A}_{27},\mathcal{A}_{27}\right)$, we obtain the algebras $\mathcal{P}_{4,62}^{\alpha},$ $\mathcal{P}_{4,63},$ $\mathcal{P}_{4,64},\mathcal{P}_{4,65},\mathcal{P}_{4,66}$.